OpenStudy (anonymous):

SAT question, how to solve? Answer given. Three children guessed the number of jelly beans in a jar. The guesses were 98, 137 and 164. None of the guesses was correct. One guess was off by 12, another by 27 and the third by 39. How many jelly beans were in the jar? ANS: 125

OpenStudy (anonymous):

164-39=125 137-12=125 98+27=125

OpenStudy (anonymous):

Yes but, how would you know which guesses were off by what amount besides trial and error...?

OpenStudy (anonymous):

@greysica is there an algebraic expression for it

OpenStudy (anonymous):

uhmm. I dont know , sorry

OpenStudy (anonymous):

Nevermind I got it x represents the amount added to the guess to make it valid pick any two guesses 164 + x = 137 + x 164 - 137 = 27 substitute into 98 + x, the third guess 98 + 27 = 125 my mathematical process is not valid at all but somehow it works

OpenStudy (anonymous):

sorry cant give any solution , but ur process is true :)